OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6>

OpenStudy (anonymous):

please help

OpenStudy (ineedhelplz):

\[v · w = v _{x} · w _{x} + v _{y} · w _{y} + vz _{2} · wz _{2} = 3 · (-4) + (-8) · (-2) + (-3) · (-6) = -12 + 16 + 18 = 22\]

OpenStudy (ineedhelplz):

It got cut off

OpenStudy (ineedhelplz):

\[-12 + 16 + 18 = 22\]

OpenStudy (anonymous):

thank you, the way you presented it really helped me understand